OpenStudy (anonymous):

L is the midpoint ofCD. If CL= 1/3x+8 and LD=2/3x-4, find the length of CD

terenzreignz (terenzreignz):

We know a midpoint cuts the line cleanly into two, right?

terenzreignz (terenzreignz):

If it's the midpoint, you can bet that the length from one endpoint to the midpoint is the same as the length of the other endpoint to the midpoint... in ENGLISH (sort of, not really) \[mCL = mLD\] \[\frac{1}{3x + 8} = \frac{2}{3x - 4}\] Can you work it out from here? :)

OpenStudy (anonymous):

si

terenzreignz (terenzreignz):

I'd be happy to know what you get as your final answer :) Anyways, peace :D